# Scalable learning of macroscopic stochastic dynamics

### v1

@@ Appendix B Theoretical Analysis
$=\tfrac{1}{K\delta t}\textstyle\sum_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}]$
@@
$=\tfrac{1}{K\delta t}\textstyle\sum_{\mathcal{I}}\mathbb{E}_{\tilde{\mathbf{x}}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}].$
Since $\mathbf{x}_{t+\delta t,\mathcal{I}}\overset{d}{=}\tilde{\mathbf{x}}_{t+\delta t,\mathcal{I}},\forall I\in\mathcal{I}$,
@@
$\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})$ $=\tfrac{1}{K\delta t}\textstyle\sum_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}]$
$=\tfrac{1}{\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}]=\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t}),$
$\boldsymbol{\Sigma}^{\ast}(\mathbf{z}_{t})$ $=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\boldsymbol{\delta}\boldsymbol{\delta}^{T}]=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\tfrac{1}{K^{2}}\textstyle\sum_{\mathcal{I},J}\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{J}}^{T}]$
@@
$=\tfrac{1}{K^{2}\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\textstyle\sum_{\mathcal{I}\neq J}\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{J}}^{T}]+\tfrac{1}{K^{2}\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\sum_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{I}}^{T}].$
Since $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}$ is independent of $\mathbf{z}_{t+\delta t,\mathcal{J}}-\mathbf{z}_{t}$ when $\mathcal{I}\neq\mathcal{J}$, we have $\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{J}}\big]$ when $\mathcal{I}\neq\mathcal{J}$. Then:

### v2

@@ 2. Error Analysis
$=\lVert\tfrac{1}{\delta t}\textstyle\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q}[\Delta\mathbf{z}_{t,\mathcal{I}}]-\tfrac{1}{\delta t}\textstyle\mathbb{E}_{\mathbf{x}_{t+\delta t}\mid\mathbf{x}_{t}}[\Delta\mathbf{z}_{t}]\rVert$
@@
$\leq\lVert\tfrac{1}{\delta t}\textstyle\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q}[\Delta\mathbf{z}_{t,\mathcal{I}}]-\tfrac{1}{\delta t}\textstyle\mathbb{E}_{\mathbf{x}_{t+\delta t}\mid\mathbf{x}_{t}}[\tfrac{1}{K}\sum_{\mathcal{I}}\Delta\mathbf{z}_{t,\mathcal{I}}]\rVert$
@@
$\textstyle+\frac{1}{\delta t}\lVert\mathbb{E}_{\mathbf{x}_{t+\delta t}\mid\mathbf{x}_{t}}[\Delta\mathbf{z}_{t}-\frac{1}{K}\sum_{\mathcal{I}}\Delta\mathbf{z}_{t,\mathcal{I}}]\rVert$
@@
$=\lVert\tfrac{1}{\delta t}\textstyle\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q}[\Delta\mathbf{z}_{t,\mathcal{I}}]-\tfrac{1}{\delta t}\textstyle\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\hat{q}}[\Delta\mathbf{z}_{t,\mathcal{I}}]\rVert$